Hi,
I'm really stuck on the math word problem on #3, #4. Can someone help me out on both of these problems on Algebra?
Thanks,
Law1736
Here is the sample question:
3. When getting into shape by exercising, the subject's maximum recommended number of heartbeats per minute (h) can be determined by subtracting the subject's age (a) from 220 and then taking 75% of that value. This relation is expressed by which of the following formulas?
4. An airplane flew for 8 hours at an airspeed of x miles per hour (mph), and for 7 more hours at 325 mph. If the average airspeed for the entire flight was 350 mph, which of the following equations could be used to find x?

Answer by rfer(16322)

3)
(220-a)(0.75)=
4)
(8x+7*325)/15=350
8x+2275=350*15
8x+2275=5250
8x=2975
x=372 mph first part